13 July 2010

On track

An icecore from 2444 m depth
The core barrel with a freshly drilled core from 2444 m depth is pulled out of the drill. It appears that the ice crystals are smaller than they’ve been over the last weeks.

After the drill got stuck on Sunday and part of Monday was used to filter the borehole, it is nice to see freshly drilled ice core in the drill trench again today. Most core breaks require rather hard pulling in the cable, and the cores are somewhat shorter than a couple of weeks ago, but with an estimated 100 m to bedrock, every drilled meter counts.

The processing line is still struggling with the Holocene brittle ice but it is planned to swap to process deep ice core again later this week. We are all excited to see what is hiding in the deep and unexplored ice. The CFA team is at full speed and has broken its own record with processing of 24.2 m ice core two days in a row.

At surface, a new 4 m deep outdoor toilet has been dug, so it seems that nothing can stop us now.

Drilling: 7.31 m, drillers depth: 2428.26 m
Logging: 6.52 m, loggers depth: 2444.25 m
Processing: bag 1451-1488 (22.0 m) 818.40 m
CFA: bag 3770-3813 (24.2m), cfa depth 2121.35 m

What we have done today:
1. Drilling and core logging ice.
2. Processed Holocene brittle ice cores.
3. CFA measurements.
4. Extended outdoor toilets.

Ad 1. Drillers report July 13
Once in a while we had an annoying problem leaving the core at bottom.
Adding the third core dog last week solved the problem, the core breaks seemed only to increase a little.
Our main, and increasing problem has been to lose contact suddenly after normal drilling. Chips collected under one of the shoes stop penetration, having drilled from 0.7 to 2.5 m core. By readjusting the drill head on barrel 2 the pitch increased from 2.7 mm to 3.3 mm. This gave several long cores. However, a new problem developed slowly, the rotation stopped suddenly at current limit, 13.2 amps. This seems to be caused by water generated at the drill head in the estimated -5°C  ice.
The core in one of these odd runs could not be broken with 2400 kg, 700 kg over a normal break. The winch controller could not hold the tension when we gave slack and the brake was automatically activated generating blue smoke. No error codes were seen on the little screen. 2.3 kg of frozen glycol was dumped in the hole, the drill was free after 2 hours. The pitch was thus reduced back to 2.7 mm, to prevent this to happen again.
Unfortunately the drilling went back to the old mode loosing contact after drilling a short core.
Our next issue is to use only one shoe, which seems not to collect chips and prevent penetration. With this configuration we were able to restart drilling when loosing contact (due to very high pitch). This has never been possible after a short run.

Weather: Mostly overcast, wind has increased to 14 knots from SE and temperatures have been -10°C  to -7°C . Pressure is dropping.

FL’s Dorthe Dahl-Jensen and Anders Svensson


← Previous entry   Next entry →